OpenStudy (anonymous):

You are rolling two dice at the same time. What is the probability of rolling a sum of 5 or 9?

OpenStudy (anonymous):

do you know either of the probabilities? if so, you add them together

OpenStudy (anonymous):

yes i do ..im just making sure i got them right

OpenStudy (anonymous):

it is easy enough to remember, but we can simply list the ways to roll a total of five (1,4), (2,3), (3,2), (4,1) so there are 4 ways out of a total of 36 and therefore the probability you get a 5 is \[\frac{4}{36}\]

OpenStudy (anonymous):

possible ways of getting a 5 are 2 and 3, 1 and 4, 3 and 2 and 4 and 1.

OpenStudy (anonymous):

similarly there are 4 ways to roll a 9, namely (3,6), (4,5), (5,4), (6,3) and again the probability is \[\frac{4}{36}\] now since you cannot roll both a 9 AND a 5, the intersection is empty, and so your OR statement becomes an addition \[\frac{4}{36}+\frac{4}{36}\]

OpenStudy (anonymous):

oh ok i get it